Circular No. 6937

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1748–288 R. M. Hjellming, M. P. Rupen, and F. Ghigo, National Radio Astron-

omy Observatory (NRAO); E. B. Waltman, Naval Research Laboratory; and A. J. Mioduszewski, Joint Institute for VLBI in Europe and NRAO, report: “Very Large Array (VLA) and Green Bank Interferometer (GBI) observations of the candidate for the radio counterpart of XTE J1748–288 (IAUC 6934) show that the radio fluxes on June 10 were factors of 3–4 above what was seen on June 7. The GBI showed preliminary fluxes at 2.25 and 8.3 GHz of ∼ 110 and ∼ 50 mJy on June 10.19 UT. At June 10.33, the VLA showed preliminary fluxes of 160, 113, 90, 66, and 55 mJy at 1.46, 4.86, 8.46, 14.9, and 22.5 GHz, respectively. This strong and highly-variable radio behavior establishes that the source reported on IAUC 6934 is the radio counterpart of the x-ray transient XTE J1748–288.”

R. P. Fender and B. W. Stappers, Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam; and M. Wieringa and R. Wark, Australia Telescope Compact Array (ATCA) Alert Team, report: “We have observed the proposed radio candidate to the x-ray transient XTE J1748– 288 (IAUC 6932, 6934) with ATCA between June 10.4 and 10.7 UT. We measure peak flux densities of 115 ± 10 and 100 ± 10 mJy at 4.8 and 8.6 GHz, respectively. The measurement at 4.8 GHz represents an increase in flux density at this frequency by a factor of four in three days (IAUC 6934), thus strengthening the association of the radio source with the x-ray transient. If the association is confirmed, this is the brightest new radio transient since the superluminal jet source GRO J1655–40, and further observations at all wavelengths are encouraged.”

SUPERNOVA 1998aq IN NGC 3982 CCD V magnitudes by D. Hanžl, Brno, Czech Republic (0.40-m re-

flector; comparison stars from GSC): May 20.912 UT, 13.1; 28.915, 13.5; 29.940, 13.56; 31.947, 13.65; June 2.921, 13.67; 7.908, 14.0; 9.906, 13.90.

SUPERNOVA 1998bu IN NGC 3368 Additional CCD photometry by Hanžl (cf. IAUC 6921): May 28.844

UT, V = 12.28 ± 0.01, B − V = +0.64 ± 0.02; 29.862, 12.38 ± 0.01, +0.71 ± 0.09; 31.873, 12.51 ± 0.01, +0.75 ± 0.02; June 2.872, 12.63 ± 0.02, +0.73 ± 0.03; 7.870, 12.84 ± 0.02, –; 9.860, 12.94 ± 0.01, +1.2 ± 0.2.

1998 June 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green